# Bongard-RWR+: Real-World Representations of Fine-Grained Concepts in Bongard Problems

### v1

@@ 1 Introduction
Abstract visual reasoning (AVR) domain stabinger2021evaluating ; van2021much refers to visual tasks, solving which requires identifying and reasoning about abstract patterns expressed through image-based analogies. Classical AVR tasks and associated benchmark problems include Raven’s Progressive Matrices (RPMs) barrett2018measuring ; zhang2019raven , visual analogy problems hill2019learning ; webb2020learning , and more malkinski2023review . These tasks mainly focus on testing the ability of systematic reasoning and generalisation to new feature distributions. Nevertheless, they typically require supervised training on large-scale datasets. This stands in contrast to human intelligence assessments, which focus on rapid adaptation to novel, potentially never-encountered problems based on prior knowledge.
@@
A distinct alternative to conventional AVR benchmarks is offered by Bongard Problems (BPs), originally proposed in 1970 bongard1970pattern . Each BP consists of two sides, each containing six images, separated according to an abstract rule. The solver’s task is to infer this underlying concept and articulate it in natural language (see Fig. 13). An alternative BP formulation relies on classifying a novel test image (or a pair of test images) to appropriate side(s). Despite its deceptively simple format, the BP setting poses several unique challenges. First, it naturally constitutes a few-shot learning problem fei2006one ; wang2020generalizing – identifying the separating concept requires generalization from just six examples per side. Second, the interpretation of images is inherently contextual linhares2000glimpse . For instance, in Fig. 1(a), a visual feature such as curvature may appear salient on the left side, but only through comparison with the opposite side the true rule, a difference in arrow directions, becomes evident. Third, solving BPs requires integrating visual perception with text generation, demanding bi-modal reasoning capabilities.
@@
Early BPs were created manually, resulting in only a few hundred instances contributed by a number of individuals foundalis2021index . To scale beyond this limitation, Bongard-LOGO nie2020bongard introduced a procedural generation method based on the action-oriented LOGO language, enabling the creation of a large set of problems. While this approach provided sufficient data for training deep models, the resulting BPs were restricted to synthetic black-and-white drawings. Subsequent efforts have shifted BPs into the real-world domain. Bongard HOI jiang2022bongard utilized natural images of human-object interactions. Bongard-OpenWorld wu2024bongardopenworld employed an open-vocabulary of free-form concepts to capture the complexity and ambiguity of real-world scenes. Most recently, Bongard-RWR malkinski2024bongardrwr ; malkinski2024reasoning linked synthetic and real-world domains by using real images to represent abstract concepts found in synthetic BPs, facilitating direct comparison of model performance across both these domains. Notably, malkinski2024reasoning  observed that Bongard-RWR, which focuses on abstract concepts, poses a greater challenge to contemporary models than Bongard HOI and Bongard-OpenWorld, which involve real-world concepts, despite all three using real-world images (see Fig. 2 for an illustration). However, since Bongard-RWR was constructed manually, its small scale limits the robustness and breadth of possible evaluations, highlighting the need for a more scalable approach to dataset construction.
@@
To overcome the scalability limitations of Bongard-RWR, we introduce Bongard-RWR+, a new benchmark featuring real-world representations of selected synthetic BPs (see an overview in Table 1). Unlike Bongard-RWR, which was constructed manually, our approach leverages recent advances in vision language models (VLMs), including Image-to-Text (I2T) and Text-to-Image (T2I) models, to automate dataset creation. For each BP in the original Bongard-RWR, we (1) use Pixtral-12B mistral2024pixtral to describe each image, (2) generate new image descriptions aligned with the matrix concept, (3) employ Flux.1-dev flux2024 to synthesize images from these descriptions, and (4) manually verify that generated images reflect the intended concept (see Fig. 3). This pipeline enables construction of a dataset comprising 5 4005\,400 BPs with the original abstract concepts preserved.
@@
  1. 1.
    We develop an automated pipeline for generating real-world-like images that express abstract visual concepts.
  2. 2.
    We introduce Bongard-RWR+, a new BP benchmark comprising 5 4005\,400 matrices constructed using this pipeline.
  3. 3.
    We conduct extensive evaluations on the proposed dataset, demonstrating that current VLMs struggle to identify fine-grained visual concepts, revealing important limitations in their AVR abilities.
@@
Figure 1: Bongard Problems. All matrices present the same abstract concept: Left side: Arrows pointing in different directions. Right side: Arrows pointing in the same direction. (a) A manually-designed synthetic BP bongard1970pattern ; foundalis2021index . (b) A manually-designed real-world representation from Bongard-RWR malkinski2024bongardrwr ; malkinski2024reasoning . (c) An automatically generated real-world representation from Bongard-RWR+.
(a) Bongard-LOGO nie2020bongard
(b) Bongard HOI jiang2022bongard
(c) Bongard-OpenWorld wu2024bongardopenworld
Figure 2: Related BP datasets. (a) A synthetic BP from Bongard-LOGO nie2020bongard ; Left: Shapes are the same. Right: Shapes are different. (b) A real-world BP from Bongard HOI jiang2022bongard ; Left: A person driving a car. Right: Not a person driving a car. (c) A real-world BP from Bongard-OpenWorld wu2024bongardopenworld ; Left: The top of a snow-covered mountain. Right: Not the top of a snow-covered mountain. Unlike Bongard-LOGO, which involves synthetic images unfamiliar to VLMs, or Bongard HOI and Bongard-OpenWorld, which focus on coarse-grained concepts, Bongard-RWR+ is designed around abstract concepts expressed through realistic images that require fine-grained visual reasoning.
@@
Synthetic BPs bongard1970pattern Synthetic Abstract $394$ $388$
Bongard-LOGO nie2020bongard Synthetic Abstract 12 00012\,000 $627$
Bongard HOI jiang2022bongard Real-world Real-world 53 00053\,000 $242$
Bongard-OpenWorld wu2024bongardopenworld Real-world Real-world 1 0101\,010 1 0101\,010
Bongard-RWR malkinski2024reasoning Real-world Abstract $60$ $55$
@@
2 Related Work
Overview of AVR challenges.
Analogy-making is a cornerstone of human cognition, closely tied to fluid intelligence—the ability to generalize learned skills to novel situations lake2017building . In the visual domain, such tasks present demanding challenges for representation learning chalmers1992high and highlight the need to integrate perception and cognition hofstadter1995fluid , motivating the development of a wide range of AVR benchmarks. Among these, RPMs raven1936mental ; raven1998raven have become a heavily studied problem type hoshen2017iq ; malkinski2024one ; malkinski2022deep ; mandziuk2019deepiq , typically involving a small set of logic-based rules. Datasets like SVRT fleuret2011comparing and those based on BPs jiang2022bongard ; malkinski2024reasoning ; nie2020bongard ; wu2024bongardopenworld introduce more abstract and diverse rules expressed through visual analogies in a few-shot learning framework. Another line of work has proposed generative challenges, such as ARC chollet2019measure and PQA qi2021pqa . Several benchmarks bitton2023vasr ; ichien2021visual ; teney2020v construct AVR tasks using real-world images, enabling evaluation of models pre-trained on large visual datasets. Our proposed dataset builds on this landscape by combining few-shot learning with real-world-like images and supporting both classification and free-form text generation tasks, offering a diverse and challenging testbed for evaluating AVR models.
BP solution methods.
In response to the wide range of AVR challenges, various problem-solving approaches have been developed hernandez2016computer ; malkinski2020multi ; mitchell2021abstraction . Specifically, for BPs, explored methods include program synthesis and inductive logic programming saito1996concept ; sonwane2021using , cognitive architectures foundalis2006phaeaco , Bayesian inference depeweg2018solving ; depeweg2024solving , convolutional networks kharagorgiev2018solving , and meta-learning approaches such as prototypical networks snell2017prototypical , SNAIL mishra2018a , or Meta-Baseline chen2021meta . More recently, LLM-based approaches have emerged, either combining image captioning models with LLMs to process text descriptions wu2024bongardopenworld or leveraging VLMs that inherently handle both modalities malkinski2024reasoning . To establish strong baselines, we conduct experiments using state-of-the-art VLMs. Additionally, prior work has framed BPs in various settings, such as binary classification where test images are assigned to matrix sides nie2020bongard ; kharagorgiev2018solving , or free-form text generation where the model articulates the underlying concept malkinski2024reasoning ; wu2024bongardopenworld . To systematically evaluate model capabilities, we cover a broad range of setups, including image (or description) to side classification, a multiclass concept selection task, and free-form text generation.
Data generation with T2I models.
Generative T2I models have gained significant attention for their broad application. Prior work has applied T2I models to generate synthetic data for supervised learning fan2024scaling ; ravuri2019classification ; sariyildiz2023fake , contrastive representation learning jahanian2022generative ; tian2024learning ; tian2023stablerep , and data augmentation trabucco2024effective . In contrast, we employ a pre-trained T2I model to automatically generate real-world-like images depicting predefined abstract concepts, enabling scalable construction of Bongard-RWR+ instances.
3 Methods
@@
Figure 3: Generative pipeline. Starting from a Bongard problem $\mathcal{BP}$ with concept $(C_{L},C_{R})$, the pipeline: (1) describes each image using an I2T model to produce paired positive and negative textual captions $\mathcal{L}^{+}_{i}$ and $\mathcal{L}^{-}_{i}$; (2) augments each positive caption with a T2T model into NN diverse descriptions $\{\mathcal{L}^{+}_{i,j}\}_{j=1}^{N}$ that preserve the underlying concept; (3) generates candidate images for each new description using a T2I model; and (4) involves a human judge to review and filter the generated images. For readability, the figure illustrates the processing flow for the first image from the left matrix side.
3.1 Bongard-RWR+
@@
To generate Bongard-RWR+, we considered each $\cal BP$} from Bongard-RWR. For each image $L_{i}\in L$ and $R_{i}\in R$, we generated paired positive and negative textual descriptions based on the side’s concept: $\text{Describe}(L_{i},C_{L})=(\mathcal{L}^{+}_{i},\mathcal{L}^{-}_{i})$ and $\text{Describe}(R_{i},C_{R})=(\mathcal{R}^{+}_{i},\mathcal{R}^{-}_{i})$, using Pixtral-12B mistral2024pixtral . This I2T model was prompted to produce positive descriptions that faithfully captured the image’s content and negative descriptions designed to steer the T2I model away from depicting the opposite concept. Each positive description was augmented with a Text-to-Text (T2T) model into $N=15$ alternative descriptions reflecting the side’s concept: $\text{Augment}(\mathcal{L}^{+}_{i},C_{L})=\{\mathcal{L}^{+}_{i,j}\}_{j=1}^{N}$ and $\text{Augment}(\mathcal{R}^{+}_{i},C_{R})=\{\mathcal{R}^{+}_{i,j}\}_{j=1}^{N}$. Each generated alternative description, paired with its corresponding negative prompt, was passed to the Flux.1-dev model flux2024 to render $512\times 512$ candidate images: $\text{Render}(\mathcal{L}^{+}_{i,j},\mathcal{L}^{-}_{i})=L_{i,j}$ and $\text{Render}(\mathcal{R}^{+}_{i,j},\mathcal{R}^{-}_{i})=R_{i,j}$. Appendix F describes the employed model prompts. All generated images were manually reviewed to ensure they accurately reflected the intended concept ($C_{L}$ or $C_{R}$) without introducing elements from the opposite side’s concept. Images failing this criterion were discarded.
@@
From the candidate images $\{L^{+}_{i,j}\}$ and $\{R^{+}_{i,j}\}$, we composed $M=10$ left and right sides, denoted $\{L^{\prime}\}_{m=1}^{M}$ and $\{R^{\prime}\}_{n=1}^{M}$, by iteratively selecting subsets that maximized intra-set visual diversity. Each subset $L^{\prime}_{m}\subset\{L^{+}_{i,j}\}$ and $R^{\prime}_{n}\subset\{R^{+}_{i,j}\}$ contained $7$ images ($6$ context images and $1$ test image), chosen iteratively to minimize total pairwise cosine similarity of ViT-L/14 embeddings dosovitskiy2021an . To cover more images and increase dataset diversity, in each iteration we selected a random image from each subset and removed it from the overall image pool in the next iterations. Finally, each left side $L^{\prime}_{m}$ was paired with each right side $R^{\prime}_{n}$ to construct $M^{2}=100$ new BP instances aligned with the original concept. Algorithm details are explained in Appendix A. Using this pipeline, we constructed 5 4005\,400 BPs corresponding to $54$ Bongard-RWR matrices ($100$ new problems per source BP). The $6$ remaining Bongard-RWR matrices were discarded due to difficulties in generating a sufficient number of images faithfully depicting the underlying concepts.
@@
We additionally introduce several dataset variants to support ablation studies. The first variant, Bongard-RWR+/GS, consists of grayscale matrices. Since Bongard-RWR+ concepts are primarily structural, as they originate from black-and-white synthetic BPs, this variant isolates the role of color, which is expected to be non-essential for concept detection. The second variant, Bongard-RWR+/LP ($P=2,\ldots,6$), e.g., denoted Bongard-RWR+/L2 for $P=2$, modifies the subset construction method by omitting the image removal step. In effect, it contains fewer unique images (1 5031\,503 in Bongard-RWR+/L6 vs. 4 1574\,157 in Bongard-RWR+), but exhibits lower average intra-side embedding similarity. Lower similarity implies greater visual diversity within a matrix, meaning the same concept is illustrated through more varied content—for example, the concept "Vertical" may be instantiated as a tree, a building, or a standing figure—potentially making the concept easier to identify. In contrast, high similarity may reflect repeated visual aspects (e.g., several images of trees), which may hint at unrelated concepts like "Nature" or "Green", making the intended concept harder to capture. We consider Bongard-RWR+/LP variants with varying numbers of images per side ($P$), enabling analysis of how the number of demonstrations influences model performance. All variants are derived from the same image pool as the main dataset. Additional details are provided in Appendix A.

### v2

@@ 1 Introduction
Abstract visual reasoning (AVR) domain (Stabinger et al., 2021; van der Maas et al., 2021) refers to visual tasks, solving which requires identifying and reasoning about abstract patterns expressed through image-based analogies. Classical AVR tasks and associated benchmark problems include Raven’s Progressive Matrices (RPMs) (Barrett et al., 2018; Zhang et al., 2019), visual analogy problems (Hill et al., 2019; Webb et al., 2020), and more (Małkiński and Mańdziuk, 2023). These tasks mainly focus on testing the ability of systematic reasoning and generalisation to new feature distributions. Nevertheless, they typically require supervised training on large-scale datasets. This stands in contrast to human intelligence assessments, which focus on rapid adaptation to novel, potentially never-encountered problems based on prior knowledge.
@@
A distinct alternative to conventional AVR benchmarks is offered by Bongard Problems (BPs), originally proposed in 1970 (Bongard, 1970). Each BP consists of two sides, each containing six images, separated according to an abstract rule. The solver’s task is to infer this underlying concept and articulate it in natural language (see Fig. 1). An alternative BP formulation relies on classifying a novel test image (or a pair of test images) to appropriate side(s). Despite its deceptively simple format, the BP setting poses several unique challenges. First, it naturally constitutes a few-shot learning problem (Fei-Fei et al., 2006; Wang et al., 2020b) – identifying the separating concept requires generalization from just six examples per side. Second, the interpretation of images is inherently contextual (Linhares, 2000). For instance, in Fig. 1(a), a visual feature such as curvature may appear salient on the left side, but only through comparison with the opposite side the true rule, a difference in arrow directions, becomes evident. Third, solving BPs requires integrating visual perception with text generation, demanding bi-modal reasoning capabilities.
@@
Early BPs were created manually, resulting in only a few hundred instances contributed by individuals (Foundalis, 2006a). To scale beyond this limitation, Bongard-LOGO (Nie et al., 2020) introduced a procedural generation method based on the action-oriented LOGO language, enabling creation of a large set of problems. While this approach provided sufficient data for training deep models, the resulting BPs were restricted to synthetic black-and-white drawings. Subsequent efforts have shifted BPs into the real-world domain. Bongard HOI (Jiang et al., 2022) utilized natural images of human-object interactions. Bongard-OpenWorld (Wu et al., 2024) employed an open-vocabulary of free-form concepts to capture the complexity and ambiguity of real-world scenes. Most recently, Bongard-RWR (Małkiński et al., 2025) linked synthetic and real-world domains by using real images to represent abstract concepts found in synthetic BPs, facilitating direct comparison of model performance across both domains. Bongard-RWR, which focuses on abstract concepts, poses a greater challenge to contemporary models than Bongard HOI and Bongard-OpenWorld, which involve real-world concepts, despite all three using real-world images (see Fig. 2 for an illustration). However, since Bongard-RWR was constructed manually, its small scale limits the robustness and breadth of possible evaluations, highlighting the need for a more scalable approach to dataset construction.
@@
To overcome the scalability limitations of Bongard-RWR, we introduce Bongard-RWR+, a new benchmark featuring real-world representations of selected synthetic BPs (see an overview in Table 1). Unlike Bongard-RWR, which was constructed manually, our approach leverages recent advances in vision language models (VLMs), including Image-to-Text (I2T) and Text-to-Image (T2I) models, to automate dataset creation. For each BP in the original Bongard-RWR, we (1) use Pixtral-12B (MistralAI, 2024) to describe each image, (2) generate new image descriptions aligned with the matrix concept, (3) employ Flux.1-dev (Labs, 2024) to synthesize images from these descriptions, and (4) manually verify that generated images reflect the intended concept (see Fig. 3). This pipeline enables construction of a dataset comprising 5 4005\,400 BPs with the original abstract concepts preserved.
@@
1) We develop a semi-automated pipeline to generate real-world-like images of abstract concepts.
2) We introduce Bongard-RWR+, a new BP benchmark of 5 4005\,400 matrices generated with this pipeline.
3) We conduct extensive evaluations on the proposed dataset, demonstrating that current VLMs struggle to identify fine-grained visual concepts, revealing important limitations in their multi-image AVR abilities.
(a) Synthetic BP #52
(b) Bongard-RWR
@@
Figure 1: Bongard Problems. All matrices present the same abstract concept: Left side: Arrows pointing in different directions. Right side: Arrows pointing in the same direction. (a) A manually-designed synthetic BP (Bongard, 1970; Foundalis, 2006a). (b) A manually-designed real-world representation from Bongard-RWR (Małkiński et al., 2025). (c) An automatically generated real-world representation from Bongard-RWR+.
(a) Bongard-LOGO
(b) Bongard HOI
(c) Bongard-OpenWorld
Figure 2: Related BP datasets. (a) A synthetic BP from Bongard-LOGO (Nie et al., 2020); Left: Shapes are the same. Right: Shapes are different. (b) A real-world BP from Bongard HOI (Jiang et al., 2022); Left: A person driving a car. Right: Not a person driving a car. (c) A real-world BP from Bongard-OpenWorld (Wu et al., 2024); Left: The top of a snow-covered mountain. Right: Not the top of a snow-covered mountain. Unlike Bongard-LOGO, which involves synthetic images unfamiliar to VLMs, or Bongard HOI and Bongard-OpenWorld, which focus on coarse-grained concepts, Bongard-RWR+ is designed around abstract concepts expressed through realistic images that require fine-grained visual reasoning.
2 Related Work
Overview of AVR challenges.
Analogy-making is a cornerstone of human cognition, closely tied to fluid intelligence—the ability to generalize learned skills to novel situations (Lake et al., 2017). In the visual domain, such tasks present demanding challenges for representation learning (Chalmers et al., 1992) and highlight the need to integrate perception and cognition (Hofstadter, 1995), motivating diverse AVR benchmarks. Among these, RPMs (Raven, 1936; Raven and Court, 1998) are an often studied problem type  (Hoshen and Werman, 2017; Małkiński and Mańdziuk, 2024b; 2025; Mańdziuk and Żychowski, 2019), typically involving a small set of logic-based rules. Datasets like SVRT (Fleuret et al., 2011) and those based on BPs (Jiang et al., 2022; Małkiński et al., 2025; Nie et al., 2020; Wu et al., 2024) introduce more abstract and diverse rules expressed through visual analogies in a few-shot learning framework. Another line of work has proposed generative challenges, such as ARC (Chollet, 2019) and PQA (Qi et al., 2021). Several benchmarks (Bitton et al., 2023; Ichien et al., 2021; Teney et al., 2020) construct AVR tasks using real-world images, enabling evaluation of models pre-trained on large visual datasets. Our proposed dataset builds on this landscape by combining few-shot learning with real-world-like images and supporting both classification and free-form text generation tasks, offering a diverse and challenging testbed for evaluating AVR models.
BP solution methods.
In response to the wide range of AVR challenges, various approaches have been developed (Hernández-Orallo et al., 2016; Małkiński and Mańdziuk, 2024a; Mitchell, 2021). Specifically, for BPs, methods include program synthesis and inductive logic programming (Saito and Nakano, 1996; Sonwane et al., 2021), cognitive architectures (Foundalis, 2006b), Bayesian inference (Depeweg et al., 2018; 2024), convolutional networks (Kharagorgiev, 2018), and meta-learning approaches such as prototypical networks (Snell et al., 2017), SNAIL (Mishra et al., 2018), or Meta-Baseline (Chen et al., 2021). Recently, LLM-based solutions combine image captioning models with LLMs for text descriptions (Wu et al., 2024) or leverage VLMs that handle both modalities (Małkiński et al., 2025; wüst2025bongardwonderlandvisualpuzzles). To establish strong baselines, we evaluate state-of-the-art VLMs. Additionally, prior work has framed BPs in various settings, such as binary classification where test images are assigned to matrix sides (Nie et al., 2020; Kharagorgiev, 2018), or free-form text generation where the model articulates the underlying concept (Małkiński et al., 2025; Wu et al., 2024). To systematically evaluate model capabilities, we cover diverse setups, including image (or description) to side classification, a multiclass concept selection task, and free-form text generation.
Data generation with T2I models.
Generative T2I models have gained significant attention for their broad application. Prior work has applied T2I models to generate synthetic data for supervised learning (Fan et al., 2024; Ravuri and Vinyals, 2019; Sarıyıldız et al., 2023), contrastive representation learning (Jahanian et al., 2022; Tian et al., 2024; 2023), and data augmentation (Trabucco et al., 2024). In contrast, we employ a pre-trained T2I model to automatically generate real-world-like images depicting predefined abstract concepts, enabling scalable construction of Bongard-RWR+ instances.
3 Methods
3.1 Bongard-RWR+
@@
Synthetic BPs (Bongard, 1970) Synthetic Abstract $394$ $388$
Bongard-LOGO (Nie et al., 2020) Synthetic Abstract 12 00012\,000 $627$
Bongard HOI (Jiang et al., 2022) Real-world Real-world 53 00053\,000 $242$
Bongard-OpenWorld (Wu et al., 2024) Real-world Real-world 1 0101\,010 1 0101\,010
Bongard-RWR (Małkiński et al., 2025) Real-world Abstract $60$ $55$
@@
Figure 3: Generative pipeline. Starting from a Bongard problem $\mathcal{BP}$ with concept $(C_{L},C_{R})$, the pipeline: (1) describes each image using an I2T model to produce paired positive/negative captions $\mathcal{L}^{+}_{i}$ and $\mathcal{L}^{-}_{i}$; (2) augments each positive caption with a T2T model into NN diverse descriptions $\{\mathcal{L}^{+}_{i,j}\}_{j=1}^{N}$ that preserve the underlying concept; (3) generates candidate images for each new description using a T2I model; and (4) involves a human judge to review and filter the generated images. For readability, the figure illustrates the processing flow for the first image from the left matrix side.
@@
To generate Bongard-RWR+, we considered each $\cal BP$} from Bongard-RWR. For each image $L_{i}\in L$ and $R_{i}\in R$, we generated paired positive and negative textual descriptions based on the side’s concept: $\text{Describe}(L_{i},C_{L})=(\mathcal{L}^{+}_{i},\mathcal{L}^{-}_{i})$ and $\text{Describe}(R_{i},C_{R})=(\mathcal{R}^{+}_{i},\mathcal{R}^{-}_{i})$, using Pixtral-12B (MistralAI, 2024). This I2T model was prompted to produce positive descriptions that faithfully captured the image’s content and negative descriptions designed to steer the T2I model away from depicting the opposite concept. Each positive description was augmented with a Text-to-Text (T2T) model into $N=15$ alternative descriptions reflecting the side’s concept: $\text{Augment}(\mathcal{L}^{+}_{i},C_{L})=\{\mathcal{L}^{+}_{i,j}\}_{j=1}^{N}$ and $\text{Augment}(\mathcal{R}^{+}_{i},C_{R})=\{\mathcal{R}^{+}_{i,j}\}_{j=1}^{N}$. Each generated alternative description, paired with its corresponding negative prompt, was passed to the Flux.1-dev model (Labs, 2024) to render $512\times 512$ candidate images: $\text{Render}(\mathcal{L}^{+}_{i,j},\mathcal{L}^{-}_{i})=L_{i,j}$ and $\text{Render}(\mathcal{R}^{+}_{i,j},\mathcal{R}^{-}_{i})=R_{i,j}$. Appendix F describes the employed model prompts. All generated images were manually reviewed to ensure they accurately reflected the intended concept ($C_{L}$ or $C_{R}$) without introducing elements from the opposite side’s concept. Images failing this criterion were discarded. Appendix E.2 presents further details.
@@
From the candidate images $\{L^{+}_{i,j}\}$ and $\{R^{+}_{i,j}\}$, we composed $M=10$ left and right sides, denoted $\{L^{\prime}\}_{m=1}^{M}$ and $\{R^{\prime}\}_{n=1}^{M}$, by iteratively selecting subsets that maximized intra-set visual diversity. Each subset $L^{\prime}_{m}\subset\{L^{+}_{i,j}\}$ and $R^{\prime}_{n}\subset\{R^{+}_{i,j}\}$ contained $7$ images ($6$ context images and $1$ test image), chosen iteratively to minimize total pairwise cosine similarity of ViT-L/14 embeddings (Dosovitskiy et al., 2021). To cover more images and increase dataset diversity, in each iteration we selected a random image from each subset and removed it from the overall image pool in the next iterations. Finally, each left side $L^{\prime}_{m}$ was paired with each right side $R^{\prime}_{n}$ to construct $M^{2}=100$ new BP instances aligned with the original concept. Algorithm details are explained in Appendix B. Using this pipeline, we constructed 5 4005\,400 BPs corresponding to $54$ Bongard-RWR matrices ($100$ new problems per source BP). The $6$ remaining Bongard-RWR matrices were discarded due to difficulties in generating a sufficient number of images faithfully depicting the underlying concepts.
@@
We additionally introduce several dataset variants to support ablation studies. The first variant, Bongard-RWR+/GS, consists of grayscale matrices. Since Bongard-RWR+ concepts are primarily structural, as they originate from black-and-white synthetic BPs, this variant isolates the role of color, which is expected to be non-essential for concept detection. The second variant, Bongard-RWR+/LP ($P=2,\ldots,6$), e.g., denoted Bongard-RWR+/L2 for $P=2$, modifies the subset construction method by omitting the image removal step. In effect, it contains fewer unique images (1 5031\,503 in Bongard-RWR+/L6 vs. 4 1574\,157 in Bongard-RWR+), but exhibits lower average intra-side embedding similarity. Lower similarity implies greater visual diversity within a matrix, meaning the same concept is illustrated through more varied content—for example, the concept "Vertical" may be instantiated as a tree, a building, or a standing figure—potentially making the concept easier to identify. In contrast, high similarity may reflect repeated visual aspects (e.g., several images of trees), which may hint at unrelated concepts like "Nature" or "Green", making the intended concept harder to capture. We consider Bongard-RWR+/LP variants with varying numbers of images per side ($P$), enabling analysis of how the number of demonstrations influences model performance. All variants are derived from the same image pool as the main dataset. Additional details are provided in Appendix B.